# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v2

@@ Appendix A Appendix
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 13: Accuracy comparison of different model scales on Ti-SafetyBench across various knowledge domains.
Task Name GPT-3.5-turbo GPT4O Claude-3.5-Sonnet Gemini-1.5-flash Deepseek-v3 Llama3.1-8B-Instruct Llama3.1-70B-Instruct
agronomy 2.96 15.38 31.95 30.18 33.14 5.33 15.98
anatomy 1.35 13.51 27.03 27.7 23.65 4.05 29.05
ancient_chinese 3.66 13.41 30.49 26.22 26.22 4.88 14.02
arts 5.0 16.25 33.75 36.25 35.62 7.5 21.25
astronomy 3.03 11.52 35.76 23.03 30.91 7.27 20.0
business_ethics 3.35 22.01 42.11 38.76 33.97 6.22 23.44
chinese_civil_service_exam 3.12 13.12 24.38 21.25 28.75 5.62 19.38
chinese_driving_rule 2.29 19.08 50.38 49.62 39.69 4.58 33.59
chinese_food_culture 2.94 19.12 27.94 34.56 29.41 4.41 29.41
chinese_foreign_policy 0.93 20.56 47.66 38.32 34.58 12.15 32.71
chinese_history 2.17 13.62 32.2 23.22 37.46 5.26 26.32
chinese_literature 2.45 18.63 23.04 28.43 29.41 7.35 19.61
chinese_teacher_qualification 2.23 17.32 45.81 36.87 25.14 6.15 22.91
clinical_knowledge 4.64 16.03 24.89 18.99 9.7 6.33 18.99
college_actuarial_science 4.72 6.6 18.87 17.92 25.47 8.49 17.92
college_education 6.54 18.69 42.99 42.06 39.25 11.21 27.1
college_engineering_hydrology 5.66 21.7 36.79 33.02 31.13 6.6 29.25
college_law 3.7 18.52 26.85 18.52 19.44 10.19 22.22
college_mathematics 6.67 11.43 20.95 19.05 28.57 8.57 18.1
college_medical_statistics 5.66 14.15 41.51 34.91 28.3 11.32 24.53
college_medicine 1.47 16.12 26.37 21.25 23.81 8.06 22.71
computer_science 4.41 14.71 25.49 29.41 29.41 8.33 24.51
computer_security 3.51 22.81 44.44 32.16 30.99 8.77 30.99
conceptual_physics 0.68 16.33 41.5 27.89 25.85 13.61 23.13
construction_project_management 2.16 14.39 38.13 29.5 28.06 8.63 12.95
economics 1.89 13.21 38.99 33.33 39.62 11.32 32.7
education 3.07 18.4 42.33 37.42 36.81 9.82 19.63
electrical_engineering 1.74 19.19 31.98 29.65 27.91 8.14 29.65
elementary_chinese 3.17 12.3 25.79 23.41 32.54 6.35 13.89
elementary_commonsense 6.06 16.16 39.9 32.83 35.86 6.06 24.24
elementary_information_and_technology 5.46 24.37 42.02 41.6 37.82 10.08 29.41
elementary_mathematics 7.39 13.04 36.09 34.78 26.09 6.96 17.39
ethnology 3.7 25.93 34.07 37.78 34.81 5.93 23.7
food_science 2.1 20.28 34.27 25.17 37.76 5.59 28.67
genetics 1.7 15.91 30.68 28.41 25.57 7.39 23.86
global_facts 2.01 19.46 34.9 31.54 38.26 8.05 19.46
high_school_biology 5.33 17.16 20.12 24.26 22.49 9.47 18.34
high_school_chemistry 4.55 12.12 20.45 12.88 16.67 13.64 19.7
high_school_geography 0.85 25.42 40.68 27.97 31.36 7.63 23.73
high_school_mathematics 6.71 12.2 34.15 25.0 37.2 7.32 20.12
high_school_physics 0.91 18.18 39.09 31.82 32.73 4.55 30.0
high_school_politics 4.2 20.28 30.07 23.08 30.77 6.29 21.68
human_sexuality 3.17 19.05 35.71 34.92 43.65 6.35 30.95
international_law 3.24 19.46 36.22 34.59 25.41 6.49 21.08
journalism 3.49 19.19 39.53 36.63 33.14 5.81 26.74
jurisprudence 2.92 19.71 39.17 30.41 37.71 9.25 21.65
legal_and_moral_basis 3.27 25.7 60.75 50.93 53.27 7.94 35.51
logical 3.25 15.45 38.21 28.46 40.65 5.69 26.02
machine_learning 0.82 11.48 29.51 27.87 22.13 6.56 21.31
management 4.76 20.0 40.48 39.05 33.81 6.19 24.76
marketing 4.44 20.56 45.0 43.33 39.44 7.22 31.11
marxist_theory 1.06 21.16 48.68 38.62 41.27 10.05 32.28
modern_chinese 5.17 18.97 24.14 22.41 20.69 8.62 12.93
nutrition 2.07 11.72 37.24 37.24 31.72 6.9 20.0
philosophy 3.81 19.05 44.76 38.1 42.86 7.62 28.57
professional_accounting 5.14 20.0 41.14 37.71 33.71 6.86 20.57
professional_law 1.9 17.06 27.96 19.43 29.38 5.69 17.06
professional_medicine 3.46 13.83 23.67 18.88 24.47 10.9 17.82
professional_psychology 4.31 21.98 37.07 34.05 37.93 6.03 25.0
public_relations 2.87 11.49 32.18 33.33 36.78 5.17 24.14
security_study 5.19 25.19 42.96 37.78 38.52 7.41 30.37
sociology 1.33 18.58 43.81 36.28 37.61 7.52 26.11
sports_science 4.24 23.64 33.33 33.33 31.52 7.88 24.85
traditional_chinese_medicine 3.24 17.84 27.03 28.11 24.86 8.65 17.3
virology 3.55 13.02 34.91 26.04 25.44 2.96 27.81
world_history 4.97 15.53 36.02 28.57 30.43 7.45 24.84
world_religions 5.0 13.12 45.62 31.87 39.38 5.0 21.25
Average 3.47 17.33 35.25 30.87 31.76 7.49 23.68
Table 14: Accuracy on 67 subjects of Ti-MMLU(CAA).The Category Mapping is in 18
Task Name Llama3.1-405B-Instruct Qwen2.5-7B-Instruct Qwen2.5-32B-Instruct Qwen2.5-72B-Instruct O1-mini Deepseek-r1
agronomy 26.04 12.43 18.34 10.65 8.28 28.99
anatomy 20.95 16.89 10.81 11.49 5.41 15.54
ancient_chinese 15.24 17.68 21.34 11.59 10.98 25.61
arts 23.75 13.12 26.25 13.12 6.88 29.38
astronomy 26.06 13.94 16.36 11.52 16.36 24.24
business_ethics 24.4 15.79 14.83 10.05 10.53 30.62
chinese_civil_service_exam 20.62 11.25 14.37 11.88 5.62 25.0
chinese_driving_rule 25.19 16.03 16.79 12.98 12.21 44.27
chinese_food_culture 24.26 11.76 17.65 15.44 6.62 36.76
chinese_foreign_policy 29.91 17.76 18.69 14.02 9.35 28.04
chinese_history 25.39 12.69 17.96 13.93 6.81 27.55
chinese_literature 21.57 15.2 19.61 14.71 13.73 26.47
chinese_teacher_qualification 34.08 15.64 20.11 13.41 10.06 41.34
clinical_knowledge 24.89 10.97 12.24 6.75 7.59 17.3
college_actuarial_science 16.04 5.66 17.92 12.26 8.49 24.53
college_education 39.25 14.02 24.3 15.89 14.95 32.71
college_engineering_hydrology 28.3 15.09 20.75 8.49 12.26 24.53
college_law 22.22 11.11 17.59 12.04 9.26 32.41
college_mathematics 18.1 13.33 10.48 13.33 3.81 9.52
college_medical_statistics 27.36 16.98 25.47 9.43 15.09 5.66
college_medicine 19.05 10.26 12.09 9.16 6.59 12.45
computer_science 28.92 10.78 15.69 9.8 11.27 20.1
computer_security 30.41 16.37 17.54 9.94 12.87 25.15
conceptual_physics 27.89 14.29 15.65 13.61 10.88 30.61
construction_project_management 23.74 16.55 19.42 12.23 7.91 28.06
economics 20.13 15.72 19.5 6.29 6.29 39.62
education 28.83 14.11 15.34 14.11 7.36 31.29
electrical_engineering 26.16 14.53 12.79 14.53 8.14 29.65
elementary_chinese 17.06 9.92 19.44 15.87 7.54 29.76
elementary_commonsense 28.28 14.14 20.2 9.6 8.08 36.36
elementary_information_and_technology 33.19 13.45 18.07 11.34 10.5 23.53
elementary_mathematics 28.7 15.22 22.61 14.78 6.52 44.35
ethnology 25.93 15.56 22.22 13.33 12.59 41.48
food_science 32.87 14.69 18.88 11.19 6.29 25.17
genetics 21.59 11.93 13.64 6.25 10.8 13.64
global_facts 20.81 10.74 18.79 9.4 10.07 18.79
high_school_biology 20.71 8.88 12.43 8.28 7.69 8.28
high_school_chemistry 19.7 7.58 12.12 6.06 9.09 6.06
high_school_geography 22.88 16.95 14.41 11.86 7.63 32.2
high_school_mathematics 24.39 26.83 24.39 18.9 9.15 27.44
high_school_physics 23.64 20.0 20.91 10.91 9.09 33.64
high_school_politics 26.57 15.38 18.18 10.49 6.99 26.57
human_sexuality 27.78 15.87 19.05 15.08 10.32 32.54
international_law 22.16 14.05 25.41 16.76 10.27 5.95
journalism 26.74 22.09 16.28 12.21 7.56 30.81
jurisprudence 27.25 15.09 19.22 11.44 9.49 32.36
legal_and_moral_basis 35.98 24.3 21.96 14.02 8.88 53.27
logical 30.08 12.2 23.58 12.2 10.57 26.83
machine_learning 24.59 9.84 12.3 6.56 6.56 18.03
management 24.76 12.86 20.0 7.14 14.76 39.52
marketing 32.22 13.89 24.44 17.22 12.22 37.22
marxist_theory 23.81 16.4 14.81 16.93 10.05 40.21
modern_chinese 19.83 14.66 20.69 10.34 6.03 29.31
nutrition 27.59 17.24 12.41 7.59 13.79 20.69
philosophy 27.62 14.29 20.0 10.48 13.33 1.9
professional_accounting 22.86 18.86 19.43 12.0 9.14 30.86
professional_law 19.43 10.9 16.59 10.43 10.43 25.12
professional_medicine 18.09 11.44 16.76 8.24 10.64 19.15
professional_psychology 21.12 11.21 23.28 13.79 9.05 33.19
public_relations 24.14 17.82 20.69 13.79 10.34 29.89
security_study 26.67 20.0 20.0 14.07 10.37 25.93
sociology 25.22 16.81 19.47 11.06 8.85 30.09
sports_science 34.55 12.12 25.45 13.33 11.52 29.7
traditional_chinese_medicine 21.08 12.97 16.76 13.51 12.43 19.46
virology 26.63 14.79 18.93 9.47 14.2 21.3
world_history 24.22 16.15 22.36 16.15 11.8 32.3
world_religions 28.12 16.25 18.75 11.25 5.62 40.62
Average 25.19 14.53 18.43 11.88 9.64 27.18
Table 15: Accuracy on 67 subjects of Ti-MMLU(CAA) continue
Task Name GPT-3.5-turbo GPT4O Claude-3.5-Sonnet Gemini-1.5-flash Deepseek-v3 Llama3.1-8B-Instruct Llama3.1-70B-Instruct
agronomy 0.59 13.02 31.36 30.18 30.18 4.73 15.98
anatomy 1.35 12.16 26.35 27.7 19.59 2.7 29.05
ancient_chinese 1.83 12.2 28.05 26.22 22.56 3.05 13.41
arts 4.38 15.0 33.12 36.25 30.0 5.0 21.25
astronomy 2.42 10.3 34.55 21.82 26.06 6.67 20.0
business_ethics 1.91 20.57 40.67 36.84 30.14 5.74 23.44
chinese_civil_service_exam 0.62 10.0 19.38 18.12 25.0 3.75 19.38
chinese_driving_rule 0.76 16.79 50.38 49.62 38.93 3.82 33.59
chinese_food_culture 2.21 17.65 27.21 33.82 27.94 4.41 29.41
chinese_foreign_policy 0.0 20.56 42.99 36.45 32.71 10.28 32.71
chinese_history 1.55 11.46 30.65 21.98 31.58 4.64 26.32
chinese_literature 1.47 17.65 22.55 28.43 27.45 4.9 19.61
chinese_teacher_qualification 2.23 16.76 44.13 35.75 24.02 4.47 22.91
clinical_knowledge 3.38 15.61 24.89 17.72 9.28 4.22 18.99
college_actuarial_science 2.83 4.72 16.98 16.98 22.64 7.55 17.92
college_education 3.74 18.69 42.06 41.12 36.45 8.41 27.1
college_engineering_hydrology 3.77 19.81 36.79 33.02 30.19 2.83 29.25
college_law 1.85 15.74 25.93 18.52 16.67 6.48 22.22
college_mathematics 2.86 9.52 17.14 16.19 20.0 5.71 18.1
college_medical_statistics 2.83 14.15 41.51 33.02 22.64 3.77 24.53
college_medicine 0.73 15.38 26.01 21.25 22.71 5.49 22.71
computer_science 2.45 12.75 25.0 28.43 27.45 4.9 24.51
computer_security 1.75 20.47 39.18 30.41 29.24 7.6 30.99
conceptual_physics 0.68 14.97 35.37 25.85 22.45 10.2 23.13
construction_project_management 1.44 12.95 35.97 28.06 25.9 5.04 12.95
economics 1.26 11.95 37.11 32.7 37.74 8.18 32.7
education 2.45 15.95 41.1 36.2 34.36 6.75 19.63
electrical_engineering 1.74 18.6 30.81 29.65 26.16 5.81 29.65
elementary_chinese 2.38 11.9 24.21 23.41 29.37 4.76 13.49
elementary_commonsense 3.54 15.66 39.9 32.83 32.83 3.54 24.24
elementary_information_and_technology 3.78 21.01 39.92 41.18 35.71 5.88 29.41
elementary_mathematics 4.35 12.17 31.74 33.91 24.35 4.35 17.39
ethnology 2.22 22.96 32.59 36.3 30.37 4.44 23.7
food_science 2.1 19.58 32.87 23.78 32.17 2.8 28.67
genetics 1.7 14.2 28.41 27.84 23.3 6.82 23.86
global_facts 0.67 18.12 33.56 30.87 38.26 7.38 19.46
high_school_biology 4.14 15.38 11.24 24.26 18.34 7.1 18.34
high_school_chemistry 0.76 9.09 13.64 10.61 14.39 10.61 19.7
high_school_geography 0.85 23.73 37.29 27.12 27.97 5.93 23.73
high_school_mathematics 5.49 11.59 31.71 25.0 33.54 5.49 20.12
high_school_physics 0.91 14.55 37.27 30.91 27.27 2.73 29.09
high_school_politics 2.1 18.18 26.57 18.88 28.67 5.59 21.68
human_sexuality 2.38 17.46 34.13 34.92 42.06 3.97 30.95
international_law 2.16 19.46 35.14 34.05 24.32 5.41 21.08
journalism 1.16 18.02 39.53 34.88 31.4 4.65 26.74
jurisprudence 1.95 17.52 35.77 27.98 36.25 8.03 21.65
legal_and_moral_basis 1.4 21.96 58.88 48.6 50.47 7.48 35.51
logical 2.44 15.45 36.59 27.64 39.02 4.07 26.02
machine_learning 0.82 10.66 23.77 27.05 20.49 4.1 21.31
management 2.86 18.1 40.0 38.1 32.86 5.71 24.76
marketing 2.22 20.0 43.89 42.78 37.22 6.11 31.11
marxist_theory 0.53 19.58 48.15 37.57 39.68 6.88 31.75
modern_chinese 3.45 18.97 20.69 21.55 18.97 3.45 12.93
nutrition 1.38 11.03 36.55 37.24 29.66 2.76 20.0
philosophy 2.86 18.1 44.76 37.14 40.0 5.71 28.57
professional_accounting 4.57 17.14 39.43 37.71 30.86 6.29 20.57
professional_law 0.95 16.11 26.07 18.96 25.12 4.74 17.06
professional_medicine 2.66 12.23 22.87 18.62 23.14 8.24 17.82
professional_psychology 2.59 18.97 36.64 34.05 35.78 4.74 25.0
public_relations 1.15 9.77 32.18 32.18 33.91 5.17 23.56
security_study 3.7 22.22 40.0 37.04 34.07 5.19 29.63
sociology 0.44 17.26 42.92 36.28 35.4 5.75 26.11
sports_science 2.42 22.42 33.33 33.33 30.3 7.27 24.85
traditional_chinese_medicine 1.62 15.68 27.03 27.57 24.32 4.86 17.3
virology 1.78 11.83 34.32 25.44 21.89 1.78 27.81
world_history 3.73 13.04 34.78 26.09 27.33 4.97 24.84
world_religions 2.5 11.88 45.62 31.87 35.0 2.5 21.25
Average 2.15 15.8 33.45 29.97 29.05 5.44 23.62
[…]

### v3

@@ C.2 Experimental Details
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 14: Accuracy Comparison of LLMs Scales on Ti-SafetyBench ($\times$100%)
Category GPT Claude Gemini DeepSeek LlaMA Qwen
3.5-turbo 4O O1-mini 3.5-Sonnet 1.5-flash v3 R1 3.1-8B 3.1-70B 3.1-405B 2.5-7B 2.5-32B 2.5-72B
business_ethics 3.35 22.01 10.53 42.11 38.76 33.97 30.62 6.22 23.44 24.40 15.79 14.83 10.05
economics 1.89 13.21 6.29 38.99 33.33 39.62 39.62 11.32 32.70 20.13 15.72 19.5 6.29
education 3.07 18.40 7.36 42.33 37.42 36.81 31.29 9.82 19.63 28.83 14.11 15.34 14.11
college_education 6.54 18.69 14.95 42.99 42.06 39.25 32.71 11.21 27.10 39.25 14.02 24.30 15.89
journalism 3.49 19.19 7.56 39.53 36.63 33.14 30.81 5.81 26.74 26.74 22.09 16.28 12.21
marketing 4.44 20.56 12.22 45.0 43.33 39.44 37.22 7.22 31.11 32.22 13.89 24.44 17.22
professional_accounting 5.14 20.00 9.14 41.14 37.71 33.71 30.86 6.86 20.57 22.86 18.86 19.43 12.00
professional_psychology 4.31 21.98 9.05 37.07 34.05 37.93 33.19 6.03 25.00 21.12 11.21 23.28 13.79
public_relations 2.87 11.49 10.34 32.18 33.33 36.78 29.89 5.17 24.14 24.14 17.82 20.69 13.79
security_study 5.19 25.19 10.37 42.96 37.78 38.52 25.93 7.41 30.37 26.67 20.00 20.00 14.07
high_school_geography 0.85 25.42 7.63 40.68 27.97 31.36 32.20 7.63 23.73 22.88 16.95 14.41 11.86
management 4.76 20.00 14.76 40.48 39.05 33.81 39.52 6.19 24.76 24.76 12.86 20.00 7.14
sociology 1.33 18.58 8.85 43.81 36.28 37.61 30.09 7.52 26.11 25.22 16.81 19.47 11.06
electrical_engineering 1.74 19.19 8.14 31.98 29.65 27.91 29.65 8.14 29.65 26.16 14.53 12.79 14.53
college_actuarial_science 4.72 6.60 8.49 18.87 17.92 25.47 24.53 8.49 17.92 16.04 5.66 17.92 12.26
college_engineering_hydrology 5.66 21.70 12.26 36.79 33.02 31.13 24.53 6.60 29.25 28.30 15.09 20.75 8.49
genetics 1.70 15.91 10.80 30.68 28.41 25.57 13.64 7.39 23.86 21.59 11.93 13.64 6.25
astronomy 3.03 11.52 16.36 35.76 23.03 30.91 24.24 7.27 20.00 26.06 13.94 16.36 11.52
college_mathematics 6.67 11.43 3.81 20.95 19.05 28.57 9.52 8.57 18.10 18.10 13.33 10.48 13.33
college_medical_statistics 5.66 14.15 15.09 41.51 34.91 28.30 5.66 11.32 24.53 27.36 16.98 25.47 9.43
virology 3.55 13.02 14.20 34.91 26.04 25.44 21.30 2.96 27.81 26.63 14.79 18.93 9.47
computer_science 4.41 14.71 11.27 25.49 29.41 29.41 20.10 8.33 24.51 28.92 10.78 15.69 9.80
conceptual_physics 0.68 16.33 10.88 41.50 27.89 25.85 30.61 13.61 23.13 27.89 14.29 15.65 13.61
anatomy 1.35 13.51 5.41 27.03 27.70 23.65 15.54 4.05 29.05 20.95 16.89 10.81 11.49
machine_learning 0.82 11.48 6.56 29.51 27.87 22.13 18.03 6.56 21.31 24.59 9.84 12.30 6.56
high_school_biology 5.33 17.16 7.69 20.12 24.26 22.49 8.28 9.47 18.34 20.71 8.88 12.43 8.28
high_school_chemistry 4.55 12.12 9.09 20.45 12.88 16.67 6.06 13.64 19.70 19.70 7.58 12.12 6.06
high_school_mathematics 6.71 12.20 9.15 34.15 25.00 37.20 27.44 7.32 20.12 24.39 26.83 24.39 18.90
high_school_physics 0.91 18.18 9.09 39.09 31.82 32.73 33.64 4.55 30.00 23.64 20.00 20.91 10.91
elementary_mathematics 7.39 13.04 6.52 36.09 34.78 26.09 44.35 6.96 17.39 28.70 15.22 22.61 14.78
legal_and_moral_basis 3.27 25.7 8.88 60.75 50.93 53.27 53.27 7.94 35.51 35.98 24.30 21.96 14.02
computer_security 3.51 22.81 12.87 44.44 32.16 30.99 25.15 8.77 30.99 30.41 16.37 17.54 9.94
food_science 2.10 20.28 6.29 34.27 25.17 37.76 25.17 5.59 28.67 32.87 14.69 18.88 11.19
college_medicine 1.47 16.12 6.59 26.37 21.25 23.81 12.45 8.06 22.71 19.05 10.26 12.09 9.16
clinical_knowledge 4.64 16.03 7.59 24.89 18.99 9.70 17.3 6.33 18.99 24.89 10.97 12.24 6.75
professional_medicine 3.46 13.83 10.64 23.67 18.88 24.47 19.15 10.90 17.82 18.09 11.44 16.76 8.24
human_sexuality 3.17 19.05 10.32 35.71 34.92 43.65 32.54 6.35 30.95 27.78 15.87 19.05 15.08
agronomy 2.96 15.38 8.28 31.95 30.18 33.14 28.99 5.33 15.98 26.04 12.43 18.34 10.65
sports_science 4.24 23.64 11.52 33.33 33.33 31.52 29.70 7.88 24.85 34.55 12.12 25.45 13.33
nutrition 2.07 11.72 13.79 37.24 37.24 31.72 20.69 6.90 20.00 27.59 17.24 12.41 7.59
elementary_information_and_technology 5.46 24.37 10.50 42.02 41.60 37.82 23.53 10.08 29.41 33.19 13.45 18.07 11.34
marxist_theory 1.06 21.16 10.05 48.68 38.62 41.27 40.21 10.05 32.28 23.81 16.40 14.81 16.93
college_law 3.70 18.52 9.26 26.85 18.52 19.44 32.41 10.19 22.22 22.22 11.11 17.59 12.04
global_facts 2.01 19.46 10.07 34.90 31.54 38.26 18.79 8.05 19.46 20.81 10.74 18.79 9.40
international_law 3.24 19.46 10.27 36.22 34.59 25.41 5.95 6.49 21.08 22.16 14.05 25.41 16.76
jurisprudence 2.92 19.71 9.49 39.17 30.41 37.71 32.36 9.25 21.65 27.25 15.09 19.22 11.44
world_religions 5.00 13.12 5.62 45.62 31.87 39.38 40.62 5.00 21.25 28.12 16.25 18.75 11.25
logical 3.25 15.45 10.57 38.21 28.46 40.65 26.83 5.69 26.02 30.08 12.20 23.58 12.20
professional_law 1.90 17.06 10.43 27.96 19.43 29.38 25.12 5.69 17.06 19.43 10.90 16.59 10.43
philosophy 3.81 19.05 13.33 44.76 38.10 42.86 1.90 7.62 28.57 27.62 14.29 20.00 10.48
world_history 4.97 15.53 11.80 36.02 28.57 30.43 32.30 7.45 24.84 24.22 16.15 22.36 16.15
arts 5.00 16.25 6.88 33.75 36.25 35.62 29.38 7.50 21.25 23.75 13.12 26.25 13.12
ancient_chinese 3.66 13.41 10.98 30.49 26.22 26.22 25.61 4.88 14.02 15.24 17.68 21.34 11.59
chinese_civil_service_exam 3.12 13.12 5.62 24.38 21.25 28.75 25.00 5.62 19.38 20.62 11.25 14.37 11.88
chinese_driving_rule 2.29 19.08 25.19 50.38 49.62 39.69 16.03 4.58 33.59 16.79 12.98 12.21 44.27
chinese_food_culture 2.94 19.12 6.62 27.94 34.56 29.41 36.76 4.41 29.41 24.26 11.76 17.65 15.44
chinese_foreign_policy 0.93 20.56 9.35 47.66 38.32 34.58 28.04 12.15 32.71 29.91 17.76 18.69 14.02
chinese_history 2.17 13.62 6.81 32.20 23.22 37.46 27.55 5.26 26.32 25.39 12.69 17.96 13.93
chinese_literature 2.45 18.63 13.73 23.04 28.43 29.41 26.47 7.35 19.61 21.57 15.20 19.61 14.71
chinese_teacher_qualification 2.23 17.32 10.06 45.81 36.87 25.14 41.34 6.15 22.91 34.08 15.64 20.11 13.41
construction_project_management 2.16 14.39 7.91 38.13 29.50 28.06 28.06 8.63 12.95 23.74 16.55 19.42 12.23
elementary_chinese 3.17 12.30 7.54 25.79 23.41 32.54 29.76 6.35 13.89 17.06 9.92 19.44 15.87
elementary_commonsense 6.06 16.16 8.08 39.90 32.83 35.86 36.36 6.06 24.24 28.28 14.14 20.20 9.60
ethnology 3.70 25.93 12.59 34.07 37.78 34.81 41.48 5.93 23.70 25.93 15.56 22.22 13.33
high_school_politics 4.20 20.28 6.99 30.07 23.08 30.77 26.57 6.29 21.68 26.57 15.38 18.18 10.49
modern_chinese 5.17 18.97 6.03 24.14 22.41 20.69 29.31 8.62 12.93 19.83 14.66 20.69 10.34
traditional_chinese_medicine 3.24 17.84 12.43 27.03 28.11 24.86 19.46 8.65 17.30 21.08 12.97 16.76 13.51
Table 15: Accuracy on 67 Subjects of Ti-MMLU (CAA) ($\times$100%)
Category GPT Claude Gemini DeepSeek LlaMA Qwen
3.5-turbo 4O O1-mini 3.5-Sonnet 1.5-flash v3 R1 3.1-8B 3.1-70B 3.1-405B 2.5-7B 2.5-32B 2.5-72B
business_ethics 1.91 20.57 4.78 40.67 36.84 30.14 19.14 5.74 23.44 24.4 1.44 12.44 7.66
economics 1.26 11.95 2.52 37.11 32.7 37.74 18.24 8.18 32.7 18.87 3.14 13.21 3.77
education 2.45 15.95 6.13 41.1 36.2 34.36 17.79 6.75 19.63 28.83 0.61 11.66 8.59
college_education 3.74 18.69 10.28 42.06 41.12 36.45 20.56 8.41 27.1 39.25 2.8 16.82 9.35
journalism 1.16 18.02 6.40 39.53 34.88 31.4 20.35 4.65 26.74 26.74 1.74 13.37 8.72
marketing 2.22 20.0 11.67 43.89 42.78 37.22 16.67 6.11 31.11 31.11 0.56 18.89 10.0
professional_accounting 4.57 17.14 6.29 39.43 37.71 30.86 15.43 6.29 20.57 22.29 0.57 13.71 5.14
professional_psychology 2.59 18.97 6.90 36.64 34.05 35.78 16.81 4.74 25.0 21.12 0.86 18.53 6.47
public_relations 1.15 9.77 8.62 32.18 32.18 33.91 18.39 5.17 23.56 24.14 1.72 17.24 5.17
security_study 3.7 22.22 5.93 40.0 37.04 34.07 12.59 5.19 29.63 26.67 3.7 14.07 14.81
high_school_geography 0.85 23.73 5.08 37.29 27.12 27.97 20.34 5.93 23.73 22.88 1.69 11.02 7.63
management 2.86 18.1 11.43 40.0 38.1 32.86 21.43 5.71 24.76 24.29 0.48 13.81 7.14
sociology 0.44 17.26 6.19 42.92 36.28 35.4 16.81 5.75 26.11 25.22 0.88 15.93 4.42
electrical_engineering 1.74 18.6 5.23 30.81 29.65 26.16 22.09 5.81 29.65 25.58 0.58 8.14 8.14
college_actuarial_science 2.83 4.72 6.60 16.98 16.98 22.64 24.53 7.55 17.92 16.04 3.77 13.21 8.49
college_engineering_hydrology 3.77 19.81 7.55 36.79 33.02 30.19 17.92 2.83 29.25 28.3 0.94 16.98 6.6
genetics 1.7 14.2 7.39 28.41 27.84 23.3 6.25 6.82 23.86 21.59 0.57 12.5 6.25
astronomy 2.42 10.3 9.7 34.55 21.82 26.06 16.97 6.67 20.0 26.06 0.61 13.94 5.45
college_mathematics 2.86 9.52 2.86 17.14 16.19 20.0 7.62 5.71 18.1 18.1 6.67 7.62 9.52
college_medical_statistics 2.83 14.15 4.72 41.51 33.02 22.64 2.83 3.77 24.53 26.42 0.94 22.64 3.77
virology 1.78 11.83 11.83 34.32 25.44 21.89 7.69 1.78 27.81 26.04 0.0 11.24 4.14
computer_science 2.45 12.75 4.41 25.0 28.43 27.45 11.27 4.9 24.51 28.43 0.49 11.27 4.41
conceptual_physics 0.68 14.97 7.48 35.37 25.85 22.45 12.24 10.2 23.13 27.89 1.36 11.56 3.4
anatomy 1.35 12.16 4.05 26.35 27.7 19.59 8.78 2.7 29.05 20.27 1.35 8.78 3.38
machine_learning 0.82 10.66 3.28 23.77 27.05 20.49 10.66 4.1 21.31 24.59 0.82 8.2 3.28
high_school_biology 4.14 15.38 6.51 11.24 24.26 18.34 1.78 7.1 18.34 20.71 1.18 7.69 4.14
high_school_chemistry 0.76 9.09 6.82 13.64 10.61 14.39 3.03 10.61 19.7 19.7 1.52 3.79 4.55
high_school_mathematics 5.49 11.59 6.71 31.71 25.0 33.54 23.78 5.49 20.12 24.39 17.68 22.56 13.41
high_school_physics 0.91 14.55 4.55 37.27 30.91 27.27 20.91 2.73 29.09 23.64 5.45 14.55 6.36
elementary_mathematics 4.35 12.17 4.78 31.74 33.91 24.35 37.39 4.35 17.39 28.26 6.09 20.0 7.83
legal_and_moral_basis 1.4 21.96 6.54 58.88 48.6 50.47 34.11 7.48 35.51 35.51 0.93 17.76 10.28
computer_security 1.75 20.47 8.77 39.18 30.41 29.24 12.28 7.6 30.99 30.41 3.51 14.04 8.77
food_science 2.10 19.58 1.40 32.87 23.78 32.17 13.29 2.8 28.67 32.87 0.00 16.08 6.99
college_medicine 0.73 15.38 3.66 26.01 21.25 22.71 6.23 5.49 22.71 19.05 1.47 6.96 3.66
clinical_knowledge 3.38 15.61 5.91 24.89 17.72 9.28 5.49 4.22 18.99 24.89 1.27 10.55 8.44
professional_medicine 2.66 12.23 7.18 22.87 18.62 23.14 6.91 8.24 17.82 18.09 0.53 12.5 6.91
human_sexuality 2.38 17.46 6.35 34.13 34.92 42.06 16.67 3.97 30.95 27.78 0.0 16.67 10.32
agronomy 0.59 13.02 5.33 31.36 30.18 30.18 20.71 4.73 15.98 25.44 0.0 13.02 7.69
sports_science 2.42 22.42 7.88 33.33 33.33 30.30 13.33 7.27 24.85 34.55 0.61 21.21 10.91
nutrition 1.38 11.03 7.59 36.55 37.24 29.66 10.34 2.76 20.0 27.59 0.69 8.28 3.45
elementary_information_and_technology 3.78 21.01 7.98 39.92 41.18 35.71 13.87 5.88 29.41 33.19 2.1 14.29 4.20
marxist_theory 0.53 19.58 7.41 48.15 37.57 39.68 22.22 6.88 31.75 23.81 2.65 10.05 8.47
college_law 1.85 15.74 8.33 25.93 18.52 16.67 17.59 6.48 22.22 21.3 1.85 12.04 6.48
global_facts 0.67 18.12 6.71 33.56 30.87 38.26 14.09 7.38 19.46 20.81 1.34 16.11 8.05
international_law 2.16 19.46 5.95 35.14 34.05 24.32 2.70 5.41 21.08 22.16 1.62 21.08 9.73
jurisprudence 1.95 17.52 7.06 35.77 27.98 36.25 17.52 8.03 21.65 27.01 1.95 12.41 9.25
world_religions 2.5 11.88 3.12 45.62 31.87 35.0 12.2 2.5 21.25 27.5 0.62 13.12 7.5
logical 2.44 15.45 8.13 36.59 27.64 39.02 12.20 4.07 26.02 30.08 2.44 17.07 5.69
professional_law 0.95 16.11 4.74 26.07 18.96 25.12 7.58 4.74 17.06 18.48 0.0 9.95 5.69
philosophy 2.86 18.1 5.71 44.76 37.14 40.0 0.95 5.71 28.57 27.62 1.9 16.19 7.62
world_history 3.73 13.04 8.07 34.78 26.09 27.33 16.77 4.97 24.84 24.22 3.11 17.39 10.56
arts 4.38 15.0 4.38 33.12 36.25 30.00 17.50 5.00 21.25 23.75 3.12 19.38 8.75
ancient_chinese 1.83 12.2 5.49 28.05 26.22 22.56 11.59 3.05 13.41 15.24 1.83 15.85 5.49
chinese_civil_service_exam 0.62 10.0 2.5 19.38 18.12 25.0 15.62 3.75 19.38 20.62 3.75 8.75 7.5
chinese_driving_rule 0.76 16.79 7.63 50.38 49.62 38.93 19.85 3.82 33.59 24.43 0.76 11.45 6.11
chinese_food_culture 2.21 17.65 2.94 27.21 33.82 27.94 18.38 4.41 29.41 24.26 0.0 13.97 7.35
chinese_foreign_policy 0.0 20.56 6.54 42.99 36.45 32.71 18.69 10.28 32.71 28.97 2.8 12.15 7.48
chinese_history 1.55 11.46 3.72 30.65 21.98 31.58 12.38 4.64 26.32 25.08 1.55 13.62 6.19
chinese_literature 1.47 17.65 6.86 22.55 28.43 27.45 19.61 4.9 19.61 21.08 1.96 16.18 9.31
[…]